OpenStudy (anonymous):

Which of the following sets of numbers is a Pythagorean triple? 14, 40, 50 7, 24, 31 7, 20, 27 14, 48, 50

OpenStudy (campbell_st):

use pythagoras' theorem to test them... that's the reason for the question \[c^2 = a^2 + b^2\]

OpenStudy (campbell_st):

c id always the longest side of the 3... so in the 1st question does \[50^2 = 40^2 + 14^2\]